Blake Scott leaves 1/8 of his estate to the boy scouts, 2/9 to the local cancer fund, and the rest to his church. What fractional part of the estate does the church recieve?
Please enter the answer as a number without the units.

Answer by whenindoubt(13)

Blake Scott estate = 100 %


1/8 goes to the BOY SCOUTS


2/9 goes to the cancer fund


Logically 1/8 + 2/9 + X = 100%, so we must find X. Therefore we must add 1/8 and 2/9 in order to do that we must find the LCD.


In this case the LCD is 72.


1/8 = 9/72


2/9 = 16/72


9/72 + 16/72 = 25/72


Therefore 47/72 remains and will go to the Church. Since there are no common factors between 47 and 72 this answer is in simplest form.